# Intraoperative 2D/3D Registration via Spherical Similarity Learning and Inference-Time Differentiable Levenberg-Marquardt Optimization

### v1

@@ 3.4 Training and Inference
Given the challenges of 2D/3D registration data annotation and collection, we adopt a self-supervised training strategy, similar to other works [15, 18]. In each iteration, a ground truth pose $\theta_{g}t$ is randomly generated near the patient’s isocenter, and a DRR is synthesized as image $I$ using a projector $\mathcal{P}(\cdot)$. The training process is divided into two stages. For the regressor $\mathcal{R}(\cdot)$ for pose initialization, we directly compute the geodesic distance in $\mathfrak{se}(3)$ between its estimation $\theta_{ini}$ and the ground truth $\theta_{gt}$, i.e. $\mathcal{L}_{\text{geo}}^{\mathfrak{se}(3)}(\theta_{ini},\theta_{gt})\$. The subsequent training of the spherical similarity learning network is guided by Eq. 15. To further mitigate the domain gap between DRRs and real X-ray images, we apply a domain randomization strategy which is similar to that in [30].
@@ 4.1 Experimental Settings
We perform experiments on three publicly available datasets to evaluate the effectiveness of the proposed method: DeepFluoro [22], Ljubljana [48], and CTSpine1K [10]. The details of the real X-ray data preprocessing and these three datasets mentioned above can be found in supplementary material  Sec. 6 and Sec. 7 respectively. We evaluated the proposed method in two application scenarios: patient-specific and patient-agnostic ( i.e. generic 2D/3D registration). In the patient-specific experiment, we performed self-supervised training on each subject individually, followed by registration with its corresponding X-ray image during the test phase. The final performance was reported as the average across all subjects. For the patient-agnostic scenario, we partitioned the dataset into training, validation, and test sets based on subjects, using a 7:2:1 split (and 4:1:1 for DeepFluoro). The evaluation was conducted on the test set, and the results were reported accordingly. Since achieving high registration accuracy solely with learning-based methods is challenging in patient-agnostic scenarios, we incorporated an optimization-based approach [6, 5] to refine the final result, enhancing the overall alignment precision.
@@
In patient-agnostic scenarios, we compared the proposed method with open-source learning-based baselines, including ProST-m [15], ProST-t [16], SOPI [8] and CDreg [7], as well as a traditional optimization-based method that utilizes BOBYQA [49] as the optimizer and multi-scale normalized cross-correlation [18] as the similarity measure. To ensure a fair comparison, all three learning-based baselines also underwent the same refinement process as the proposed method and used the same pose initialization network. As the smaller sample sizes in the other two datasets limited their suitability for this analysis, we conducted experiments exclusively on the CTSpine1K dataset, which has the largest number of subjects. In the patient-specific scenario, our baseline methods include PSSS-reg [64], DiffPose [18], PoseNet [3], DFLNet [22] and SCR-reg [54]. We follow the original parameter settings of other methods except for an additional grid search for the hyperparameters of SCR-reg. To be note that, since the Ljubljana dataset lacks landmark annotations, landmark-based methods (DFLNet and SCR-reg) cannot be applied to this dataset. The implementation details of the proposed method is provided in supplementary material Sec. 8
@@ 4.2 Results
PSSS-reg [64] 56.0% 0.93 2.51 5.57 12.7 s
PoseNet [3] 4.3% 16.6 22.0 29.2 0.1 s
DFLNet [22] 36.6% 3.20 7.29 13.1 1.0 s
SCR-reg [54] 33.3% 4.70 9.59 12.8 1.1 s
DiffPose [18] 83.1% 0.60 0.89 1.47 5.3 s
@@
PSSS-reg [64] 40.0% 2.48 5.87 11.3 15.3 s
PoseNet [3] 0% 23.3 26.2 29.2 <<0.1 s
DiffPose [18] 80.0% 0.63 0.94 1.78 6.0 s
@@
PSSS-reg [64] 31.4% 4.57 9.75 15.8 16.2 s
PoseNet [3] 9.8% 11.7 17.2 24.5 << 0.1 s
DFLNet [22] 28.4% 4.80 10.9 18.1 1.3 s
SCR-reg [54] 22.2% 7.08 12.6 19.7 1.3 s
DiffPose [18] 66.4% 0.77 1.51 3.39 7.3 s
@@
ProST-m [15] 37.6% 3.03 7.36 13.6 18.7 s
ProST-t [16] 46.3% 2.03 9.56 20.4 13.2 s
SOPI [8] 43.8% 1.99 6.28 12.5 14.2 s
CDreg [7] 50.1% 0.99 7.72 17.4 10.1 s
@@ 6 Preprocessing of Real X-ray Data
As mentioned in numerous prior studies [18, 60, 44], X-ray imaging captures the intensity attenuation of X-rays as they traverse a medium, whereas DRRs quantify the absorption of X-ray energy by the medium. Therefore, in DRRs, higher-density bone tissue appears brighter, while lower-density soft tissue appears relatively darker. This contrast is inverted in X-ray images, where bone structures appear darker due to greater X-ray absorption, while soft tissue appears brighter due to lower attenuation. To ensure that the grayscale distribution of the DRR generated using Siddon’s method [56] aligns with that of X-ray images governed by the Beer-Lambert law, we applied the following inversion strategy:
@@ 7 Datasets
  • DeepFluoro: The DeepFluoro dataset contains six pelvic CT scans with a total of 366 real fluoroscopic X-ray images [22]. It provides the calibrated intrinsic matrix of the C-arm imaging system along with the ground truth extrinsic matrix for each X-ray image. Additionally, 14 manually annotated landmarks are available for each CT scan.
  • Ljubljana: A clinical dataset [48] contains ten clinical 3D Cone-Beam Computed Tomography (CBCT) subtracted angiography scans of patients undergoing cerebral endovascular treatments. Each CBCT scan is provided with two corresponding X-ray images. Calibrated intrinsic and extrinsic matrices are provided.
  • CTSpine1k: The CTSpine1K dataset [10] consists of 1005 CT volumes with segmentation masks collected from four different open sources, capturing diverse appearance variations. After quality control, 651 CT scans were retained. Following segmentation and denoising preprocessing, only the lumbar spine was extracted. The contour points of the spine surface were obtained by segmentation using Hounsfield unit (HU) threshold, morphological refinement, and 3D surface extraction using the Marching Cubes algorithm. The automatic landmark identification follows a strategy similar to [4, 65]. For each CT scan, 500 DRR images with random poses were generated using [60]. The intrinsic matrix used in the simulation environment is calibrated to match the settings of a mobile C-arm imaging device.
@@ 8 Implementation Details
For patient-specific 2D/3D registration, we selected ResNet-18 as the backbone for the pose regressor $\mathcal{R}(\cdot)$, using the Adam optimizer with a learning rate of 1e-3 and a cosine scheduler. The spherical similarity network, used for subsequent fine-tuning, was trained with the AdamW optimizer, also with a learning rate of 1e-3, but utilizing a cyclic scheduler. The pose regressor and similarity network were trained in a self-supervised manner using 500k and 200k synthetic images generated in real-time, respectively. For the patient-agnostic scenario, the parameter settings remained largely the same, except that the pose regressor was replaced with RTPIv3 [8], which incorporates spatial information from CT. Additionally, the number of self-supervised training iterations was increased to 750k for the pose regressor and 300k for the similarity network, to better generalize across different subjects. The termination criterion for the differentiable LM optimization is set such that the standard deviation of $\Delta\theta_{i}$ over the last ten iterations is less than 1e-2. This threshold was determined through grid search across all datasets to ensure an optimal balance between convergence stability and computational efficiency. Our experiments were conducted on a PC equipped with an NVIDIA RTX 3090 GPU and an Intel Core i7 CPU. Due to memory limitations, the batch sizes for training the pose regressor and the spherical similarity network were set to 4 and 1, respectively.
@@ 9 Domain Randomization Strategy
Similar to [20, 18, 64], to alleviate the domain gap between synthetic image and real X-ray image, domain randomization strategy is adopted during training:
@@
  • Random Erasing [68]: Randomly erases a rectangular region of the image with an area sampled from $0.02\cdot\text{area},0.4\cdot\text{area}]$ and an aspect ratio sampled from $[0.3,1/0.3]$, filling the region with the mean intensity of the image.

### v3

@@ 3.4 Training and Inference
Given the challenges of 2D/3D registration data annotation and collection, we adopt a self-supervised training strategy, similar to other works [gao2020generalizing, gopalakrishnan2024intraoperative]. In each iteration, a ground truth pose $\theta_{gt}$ is randomly generated near the patient’s isocenter, and a DRR is synthesized as image $I$ using a projector $\mathcal{P}(\cdot)$. The training process is divided into two stages. For the regressor $\mathcal{R}(\cdot)$ for pose initialization, we directly compute the geodesic distance in $\mathfrak{se}(3)$ between its estimation $\theta_{ini}$ and the ground truth $\theta_{gt}$, i.e. $\mathcal{L}_{\text{geo}}^{\mathfrak{se}(3)}(\theta_{ini},\theta_{gt})\$. The subsequent training of the spherical similarity learning network is guided by Eq. 15. To further mitigate the domain gap between DRRs and real X-ray images, we apply a domain randomization strategy which is similar to that in [jaganathan2023self].
@@ 4.1 Experimental Settings
We perform experiments on three publicly available datasets to evaluate the effectiveness of the proposed method: DeepFluoro [grupp2020automatic], Ljubljana [pernus20133d], and CTSpine1K [deng2021ctspine1k]. The details of the real X-ray data preprocessing and these three datasets mentioned above can be found in supplementary material  Sec. 6 and Sec. 7 respectively. We evaluated the proposed method in two application scenarios: patient-specific and patient-agnostic ( i.e. generic 2D/3D registration). In the patient-specific experiment, we performed self-supervised training on each subject individually, followed by registration with its corresponding X-ray image during the test phase. The final performance was reported as the average across all subjects. For the patient-agnostic scenario, we partitioned the dataset into training, validation, and test sets based on subjects, using a 7:2:1 split (and 4:1:1 for DeepFluoro). The evaluation was conducted on the test set, and the results were reported accordingly. Since achieving high registration accuracy solely with learning-based methods is challenging in patient-agnostic scenarios, we incorporated an optimization-based approach [chen2024optimization, chen2025introducing] to refine the final result, enhancing the overall alignment precision.
@@
In patient-agnostic scenarios, we compared the proposed method with open-source learning-based baselines, including ProST-m [gao2020generalizing], ProST-t [gao2023fully], SOPI [chen2024embedded] and CDreg [chen2024fully], as well as a traditional optimization-based method that utilizes BOBYQA [powell2009bobyqa] as the optimizer and multi-scale normalized cross-correlation [gopalakrishnan2024intraoperative] as the similarity measure. To ensure a fair comparison, all three learning-based baselines also underwent the same refinement process as the proposed method and used the same pose initialization network. As the smaller sample sizes in the other two datasets limited their suitability for this analysis, we conducted experiments exclusively on the CTSpine1K dataset, which has the largest number of subjects. In the patient-specific scenario, our baseline methods include PSSS-reg [zhang2023patient], DiffPose [gopalakrishnan2024intraoperative], PoseNet [bui2017x], DFLNet [grupp2020automatic] and SCR-reg [shrestha2023x]. We follow the original parameter settings of other methods except for an additional grid search for the hyperparameters of SCR-reg. To be note that, since the Ljubljana dataset lacks landmark annotations, landmark-based methods (DFLNet and SCR-reg) cannot be applied to this dataset. The implementation details of the proposed method is provided in supplementary material Sec. 8
@@ 4.2 Results
PSSS-reg [zhang2023patient] 56.0% 0.93 2.51 5.57 12.7 s
PoseNet [bui2017x] 4.3% 16.6 22.0 29.2 0.1 s
DFLNet [grupp2020automatic] 36.6% 3.20 7.29 13.1 1.0 s
SCR-reg [shrestha2023x] 33.3% 4.70 9.59 12.8 1.1 s
DiffPose [gopalakrishnan2024intraoperative] 83.1% 0.60 0.89 1.47 5.3 s
@@
PSSS-reg [zhang2023patient] 40.0% 2.48 5.87 11.3 15.3 s
PoseNet [bui2017x] 0% 23.3 26.2 29.2 <<0.1 s
DiffPose [gopalakrishnan2024intraoperative] 80.0% 0.63 0.94 1.78 6.0 s
@@
PSSS-reg [zhang2023patient] 31.4% 4.57 9.75 15.8 16.2 s
PoseNet [bui2017x] 9.8% 11.7 17.2 24.5 << 0.1 s
DFLNet [grupp2020automatic] 28.4% 4.80 10.9 18.1 1.3 s
SCR-reg [shrestha2023x] 22.2% 7.08 12.6 19.7 1.3 s
DiffPose [gopalakrishnan2024intraoperative] 66.4% 0.77 1.51 3.39 7.3 s
@@
ProST-m [gao2020generalizing] 37.6% 3.03 7.36 13.6 18.7 s
ProST-t [gao2023fully] 46.3% 2.03 9.56 20.4 13.2 s
SOPI [chen2024embedded] 43.8% 1.99 6.28 12.5 14.2 s
CDreg [chen2024fully] 50.1% 0.99 7.72 17.4 10.1 s
@@ 6 Preprocessing of Real X-ray Data
As mentioned in numerous prior studies [gopalakrishnan2024intraoperative, unberath2018deepdrr, momeni2024voxel], X-ray imaging captures the intensity attenuation of X-rays as they traverse a medium, whereas DRRs quantify the absorption of X-ray energy by the medium. Therefore, in DRRs, higher-density bone tissue appears brighter, while lower-density soft tissue appears relatively darker. This contrast is inverted in X-ray images, where bone structures appear darker due to greater X-ray absorption, while soft tissue appears brighter due to lower attenuation. To ensure that the grayscale distribution of the DRR generated using Siddon’s method [siddon1985fast] aligns with that of X-ray images governed by the Beer-Lambert law, we applied the following inversion strategy:
@@ 7 Datasets
  • DeepFluoro: The DeepFluoro dataset contains six pelvic CT scans with a total of 366 real fluoroscopic X-ray images [grupp2020automatic]. It provides the calibrated intrinsic matrix of the C-arm imaging system along with the ground truth extrinsic matrix for each X-ray image. Additionally, 14 manually annotated landmarks are available for each CT scan.
  • Ljubljana: A clinical dataset [pernus20133d] contains ten clinical 3D Cone-Beam Computed Tomography (CBCT) subtracted angiography scans of patients undergoing cerebral endovascular treatments. Each CBCT scan is provided with two corresponding X-ray images. Calibrated intrinsic and extrinsic matrices are provided.
  • CTSpine1k: The CTSpine1K dataset [deng2021ctspine1k] consists of 1005 CT volumes with segmentation masks collected from four different open sources, capturing diverse appearance variations. After quality control, 651 CT scans were retained. Following segmentation and denoising preprocessing, only the lumbar spine was extracted. The contour points of the spine surface were obtained by segmentation using Hounsfield unit (HU) threshold, morphological refinement, and 3D surface extraction using the Marching Cubes algorithm. The automatic landmark identification follows a strategy similar to [campbell2015automated, zhang2024spineclue]. For each CT scan, 500 DRR images with random poses were generated using [unberath2018deepdrr]. The intrinsic matrix used in the simulation environment is calibrated to match the settings of a mobile C-arm imaging device.
@@ 8 Implementation Details
For patient-specific 2D/3D registration, we selected ResNet-18 as the backbone for the pose regressor $\mathcal{R}(\cdot)$, using the Adam optimizer with a learning rate of 1e-3 and a cosine scheduler. The spherical similarity network, used for subsequent fine-tuning, was trained with the AdamW optimizer, also with a learning rate of 1e-3, but utilizing a cyclic scheduler. The pose regressor and similarity network were trained in a self-supervised manner using 500k and 200k synthetic images generated in real-time, respectively. For the patient-agnostic scenario, the parameter settings remained largely the same, except that the pose regressor was replaced with RTPIv3 [chen2024embedded], which incorporates spatial information from CT. Additionally, the number of self-supervised training iterations was increased to 750k for the pose regressor and 300k for the similarity network, to better generalize across different subjects. The termination criterion for the differentiable LM optimization is set such that the standard deviation of $\Delta\theta_{i}$ over the last ten iterations is less than 1e-2. This threshold was determined through grid search across all datasets to ensure an optimal balance between convergence stability and computational efficiency. Our experiments were conducted on a PC equipped with an NVIDIA RTX 3090 GPU and an Intel Core i7 CPU. Due to memory limitations, the batch sizes for training the pose regressor and the spherical similarity network were set to 4 and 1, respectively.
@@ 9 Domain Randomization Strategy
Similar to [grimm2021pose, gopalakrishnan2024intraoperative, zhang2023patient], to alleviate the domain gap between synthetic image and real X-ray image, domain randomization strategy is adopted during training:
@@
  • Random Erasing [zhong2020random]: Randomly erases a rectangular region of the image with an area uniformly sampled from $0.02\cdot\text{area},0.4\cdot\text{area}]$, and an aspect ratio sampled uniformly from $[0.3,1]$, filling the region with the mean intensity of the whole image.